# A Physics-Inspired Optimizer: Velocity Regularized Adam

### v1

@@ B.1 Image Classification
Scheduler Type WarmupCosineAnnealing
@@
VRAdam $\eta$ Log-uniform, Min: $1\times 10^{-4}$, Max: 0.1 0.0846
@@
VRAdam normgrad Values: [True, False] True
VRAdam lr cutoff Integer Uniform, Min: 3, Max: 30 29
AdamW $\eta$ Log-uniform, Min: $1\times 10^{-5}$, Max: $1\times 10^{-1}$ 0.0625
SGD $\eta$ Log-uniform, Min: $1\times 10^{-5}$, Max: $1\times 10^{-1}$ 0.00784
RMSProp $\eta$ Log-uniform, Min: $1\times 10^{-5}$, Max: 0.1 1.78e-4
B.2 Language Modeling
@@
VRAdam $\eta$ Log-uniform, Min: $1\times 10^{-5}$, Max: 0.1 1.55e-05
@@
VRAdam lr cutoff Integer Uniform, Min: 5, Max: 30 7
@@
SGD $\eta$ Log-uniform, Min: $1\times 10^{-5}$, Max: $1\times 10^{-1}$
@@
B.3 Generative Modeling with Diffusion
This section details the comprehensive sweep for the diffusion model, on the MNIST dataset.
  • Model: UNet with self-attention
  • Dataset: MNIST
  • Hyperparameter sweep method: Bayesian optimization
  • Optimization metric: validation loss
Table 10: Dataset Splits during Diffusion Comprehensive Sweep
Dataset Split Configuration Description
Training 90% of the original MNIST training set (60,000 images).
Validation 10% of the original MNIST training set.
Test Full MNIST test set (10,000 images).
Table 11: Fixed Hyperparameters during Diffusion Comprehensive Sweep
Parameter Value
Seed 42
Epochs 50
Batch Size 128
UNet base channels 96
UNet time embed dim 128
UNet use attention True
Diffusion num timesteps 400
Diffusion $\beta_{min}$ $1\times 10^{-4}$
Diffusion $\beta_{max}$ 0.02
Sample Every (sample every) 5 epochs
Optimizer weight decay (Common) $1\times 10^{-5}$
Adam/VRAdam $\beta_{1}$ 0.9
Adam/VRAdam $\beta_{2}$ 0.999
Adam/VRAdam $\epsilon$ $1\times 10^{-8}$
VRAdam power 2
SGD momentum 0.2
SGD nesterov True
RMSProp alpha 0.1
RMSProp $\epsilon$ $1\times 10^{-8}$
RMSProp momentum 0.2
Table 12: Swept Hyperparameters during Diffusion Comprehensive Sweep
Optimizer Parameter Sweep Configuration Optimal Parameter
VRAdam $\eta$ Log-uniform, Min: $1\times 10^{-5}$, Max: $1\times 10^{-1}$ $3.38\times 10^{-5}$
VRAdam $\beta_{3}$ Uniform, Min: 0.1, Max: 2.0 1.99
VRAdam lr cutoff Integer Uniform, Min: 5, Max: 30 6
VRAdam normgrad Values: [True, False] True
Adam $\eta$ Log-uniform, Min: $1\times 10^{-5}$, Max: $1\times 10^{-1}$ $3.24\times 10^{-5}$
SGD $\eta$ Log-uniform, Min: $1\times 10^{-5}$, Max: $1\times 10^{-1}$ 0.00227
RMSProp $\eta$ Log-uniform, Min: $1\times 10^{-5}$, Max: $1\times 10^{-1}$ $1.81\times 10^{-5}$
Method Validation Loss Test Loss
VRAdam 0.0360 0.0360
AdamW 0.0354 0.0354
SGD+Nesterov 0.056 0.054
RMSProp 0.0363 0.0350
Table 13: Diffusion model results on MNIST.
B.4 Generative Modeling with GFlowNets
@@
B.5 Edge of Stability analysis
@@
Appendix C Compute Resources
All experiments were run on Lambda cloud instances or the Google cloud platform (GCP). Experiments were conducted either using a NVIDIA L4 GPU with 24 GB of GPU memory and 31 GB of system memory or larger experiments were performed on a NVIDIA A10 with 24 GB of GPU memory and 200 GB of system memory.
Appendix D Loss curves with Error Envelopes
@@
Figure 12: Train (left) and validation (right) loss curves calculated using different run values for the diffusion model using SGD Nesterov with momentum.
Figure 13: Train (left) and validation (right) loss curves calculated using different run values for the diffusion model using RMSProp.

### v2

@@ D.1 Image Classification
Scheduler Type (AdamW) WarmupCosineAnnealing
@@
VRAdam $\alpha_{0}$ Log‐uniform, Min: $1\times 10^{-4}$, Max: 0.1 0.0846
@@
VRAdam $\alpha_{1}$ Integer Uniform, Min: 3, Max: 30 29
AdamW $\eta$ Log‐uniform, Min: $1\times 10^{-5}$, Max: $1\times 10^{-1}$ 0.0625
SGD $\eta$ Log‐uniform, Min: $1\times 10^{-5}$, Max: $1\times 10^{-1}$ 0.00784
RMSProp $\eta$ Log‐uniform, Min: $1\times 10^{-5}$, Max: 0.1 1.78e-4
D.2 Language Modeling
@@
VRAdam $\alpha_{0}$ Log-uniform, Min: $1\times 10^{-5}$, Max: 0.1 1.55e-05
@@
VRAdam $\alpha_{1}$ Integer Uniform, Min: 5, Max: 30 7
@@
SGD $\eta$ Log-uniform, Min: $1\times 10^{-5}$, Max: $1\times 10^{-1}$ -
@@
D.3 Generative Modeling with GFlowNets
@@
D.4 Edge of Stability analysis
@@
Appendix E Compute Resources
All experiments were run on Lambda cloud instances or the Google cloud platform (GCP). Experiments were conducted either using a NVIDIA L4 GPU with 24 GB of GPU memory and 31 GB of system memory or larger experiments were performed on a NVIDIA A10 with 24 GB of GPU memory and 200 GB of system memory. The GPT benchmark was run on an NVIDIA H100 GPU.
Appendix F Loss curves with Error Envelopes